OpenStudy (anonymous):

How could I find the graph of this equation 2x - 3y = 12 ? What would be the slope and the y-intercept ..

OpenStudy (anonymous):

Yes , I thought I knew how to convert it but I got a whole bunch of decimals . like Y = 1.3x - 2.7 ?

OpenStudy (anonymous):

So we're turning 2x-3y=12 into slope intercept form; y=mx+b to graph 2x-3y=12 We want y by itself so subtract both sides by 2x -3y=-2x+12 Now we divide both sides by -3 y=2/3x-4

OpenStudy (anonymous):

Do you understand that?

OpenStudy (anonymous):

Yes I think so . So the slope would be 2/3 and the y intercept would be at -4 ?

OpenStudy (anonymous):

Exactly. Did you need help graphing, or do you got it from here?

OpenStudy (anonymous):

These are the two choices I cannot chose between . They look the same to me .

OpenStudy (anonymous):

hmm....They look like the same exact graph to me.

OpenStudy (anonymous):

They look the same to me as well . maybe it was a mistake in the test . Thank you for the help !

OpenStudy (anonymous):

Those are actually both wrong

OpenStudy (anonymous):

The graph should be going the opposite way. Your slope is 2/3, so you would be going up 2 and to the right 3

OpenStudy (anonymous):

Do you have that as a choice?

OpenStudy (anonymous):

This one ?

OpenStudy (anonymous):

That would be it!